OpenStudy (anonymous):

[9.05] Given the function g(x) = 4x − 5, find g(−3).

OpenStudy (anonymous):

in place of x put -3 g(-3) = 4*(-3) -5 =-12-5 =-17

OpenStudy (shamim):

agreee with @Koel

OpenStudy (anonymous):

Perfect Thanks(: